4x+2y=-8
2x-4=-16

Answer by sachi(548)

4x+2y=-8-------------1
2x-4=-16-----------2
from eqn 1 x=-(8+2y)/4----------3
from 2 7 3 2[-(8+2y)/4]-4=-16
-(8+2y)/4=(-16+4)/2=-6
or 8+2y=6*4=24
or 2y=24-8=16
or y=16/2=8 & from 3 x=-(8+2*8)/4=-6 ans